OpenStudy (anonymous):

a farmer is planning a rectangular area for her chickens. the area of the ractangle will be 200 square feet. three sides of the rectangle will be formed by fencing which cost $5 per foot. the fourth side of the rectangle will be formed by a portion of the barn wallm which requires no fencing. in order to minimzie the cost of fencing how long should the fourth side be?

OpenStudy (anonymous):

lets get a picture 

OpenStudy (anonymous):

total cost is \(5x+5x+5y=10x+5y\) and since \(xy=200\) you have \(y=\frac{200}{x}\)

OpenStudy (anonymous):

now \[C(x)=10x+\frac{40}{x}\] and you want the minimum

OpenStudy (anonymous):

take the derivative set it equal to zero to find the critical points etc

OpenStudy (anonymous):

how

OpenStudy (anonymous):

derivative set?

OpenStudy (anonymous):

take the derivative find the critical points by "setting it equal to zero"

OpenStudy (anonymous):

that is so confusing oh ym god.

OpenStudy (anonymous):

is this a calculus course?

OpenStudy (anonymous):

geometry

OpenStudy (anonymous):

oh then in that case i have no idea what you are supposed to do

OpenStudy (anonymous):

hahahaha oh god ok thanks anyways

OpenStudy (anonymous):

i can tell you that the long side which is parallel to the side against the barn will be twice the length of the other two sides so it will be 100 and the other two will be 50 each

OpenStudy (anonymous):

but doing this using geometry and not calculus is above my level of expertise i have no idea how to do it using just geometry